# A finite element method for anisotropic crystal growth on surfaces

### v1

@@ 4 Finite element approximation
Let $\mathcal{M}^{h}$ be a polyhedral hypersurface approximating $\mathcal{M}$, and let $\{{\cal T}^{h}\}_{h>0}$ be a family of open triangles with $\mathcal{M}^{h}=\cup_{\sigma\in{\cal T}^{h}}\overline{\sigma}$. We refer to ? for more details on polyhedral approximations of surfaces and finite element spaces on polyhedral surfaces. Associated with ${\cal T}^{h}$ is the finite element space
@@ 4.1 The obstacle potential
It is possible to generalize the considered model in the following ways. All of the results presented in this paper remain valid for these generalizations, on using the techniques developed by the authors in ?.
@@
    can be considered, so that (3.4) collapses to the case $r=1$. In the case $r>1$, the term $B(\cdot,\nabla_{\!s}\Phi^{n-1})$ in, e.g., (4.2b) needs to be changed to $B_{r}(\cdot,\nabla_{\!s}\Phi^{n-1},\nabla_{\!s}\Phi^{n})$, where the matrices $B_{r}$ are defined analogously to the flat and spatially homogeneous case in ?.
@@
Proof. The existence and uniqueness results follow analogously to Lemma 3.1 in ?, and so we omit their proof here.
@@ 4.2 Smooth potentials
and for simplicity we restrict our attention to that case. Details on how to deal with more general potentials can be found in ?.
@@
Proof. The existence and uniqueness results follow analogously to Theorem 3.8 in ?. The proof of the stability bound (4.2) is similar to the proof of Theorem 4.2, making use of the convex/concave splitting $\Psi^{\prime}=\phi=\phi^{+}+\phi^{-}$.
@@ 5 Numerical results
We implemented the scheme (4.1) with the help of the finite element toolbox ALBERTA, see ?. To increase computational efficiency, we employ adaptive meshes, which have a finer mesh size $h_{f}\approx\frac{1}{N_{f}}$ within the diffuse interfacial regions and a coarser mesh size $h_{c}\approx\frac{1}{N_{c}}$ away from them, with $N_{f},N_{c}\in{\mathbb{N}}$, see ?? for a more detailed description in the planar case. For the solution of the nonlinear systems of equations we employ the iterative solvers discussed in ?.
@@ 5.1 Spatially inhomogeneous anisotropies in 2d
Then, on $\mathcal{M}=(-\frac{3}{2},\frac{3}{2})\times(\frac{1}{2},\frac{1}{2})\times\{0\}$ we use the spatially weighted isotropic surface energy
@@
Figure 2: ($\varepsilon=(16\pi)^{-1}$) Solutions for the Cahn–Hilliard equation in the case of a spatially inhomogeneous energy in the plane. Snapshots of the evolutions at times $t=0,0.1,\ldots,0.7,2$.
@@ 5.2 Spatially homogeneous anisotropies in 3d
where $R_{1}(\theta)=\left(\!\!\!\scriptsize\begin{array}[]{rrr}\cos\theta&\sin\theta&0\\ -\sin\theta&\cos\theta&0\\ 0&0&1\end{array}\!\!\right)$, $R_{2}(\theta)=\left(\!\!\!\scriptsize\begin{array}[]{rrr}\cos\theta&0&\sin\theta\\ 0&1&0\\ -\sin\theta&0&\cos\theta\end{array}\!\!\right)$ and $l_{\delta}({p})=\left[\delta^{2}\,|{p}|^{2}+p_{1}^{2}\,(1-\delta^{2})\right]^{\frac{1}{2}}$. The Wulff shapes for (5.1) are shown in Figure 3. We recall from ? that for a given anisotropy $\gamma_{0}$ the boundary of its Wulff shape is the solution of the isoperimetric problem for $\mathcal{E}_{0}(\Gamma)=\int_{\Gamma}\gamma_{0}(\nu)\;{\rm d}{\mathcal{H}}^{2}$. Hence they are often used to visualize the properties of $\gamma_{0}$.
@@
We note that for $\delta=0$, the densities (5.1) represent crystalline anisotropies, whose Wulff shapes are given by hexagonal prisms, cf. ?, Fig. 3. The coefficients in (5.1)(a) are chosen such that, in the limit $\delta=0$, the eight admissible normal directions all have the same energy, while in (5.1)(b) and (5.1)(c) the two normals ±$\pm e_{3}$ have a smaller/larger energy than the other six directions.
We use the anisotropy (5.1)(b) with $\delta=10^{-2}$ to numerically investigate the statement from Lemma 3.1. To this end, we choose as initial interface a perturbed great circle on the unit sphere, that lies in a hyperplane that makes a certain fixed angle $\theta$ with the $x-z$-plane. In particular, in Figure 4 we show simulations for $\theta=10^{\circ},30^{\circ},60^{\circ}$. Here we let $\varepsilon=(16\pi)^{-1}$, $N_{c}=1$, $N_{f}=128$ and $\tau=10^{-4}$. We note that each simulations settles on a great circle on the sphere. Here we note that as $\nu_{\min}=e_{3}$, the equator has the lowest energy of all the great circles, and so if the initial curve on $\mathcal{M}$ is sufficiently inclined, that global minimizer is indeed reached by the evolution. For only small deviations from a north-south great circle, the evolutions settles on that local minimizer instead.
Figure 4: ($\varepsilon=(16\pi)^{-1}$) Anisotropic Cahn–Hilliard equation on a sphere. Snapshots of the evolutions at times $t=0,1,5$ (from left to right).
@@ 5.2.2 Convergence experiment on the unit sphere
We use the rotationally symmetric solution for the Mullins–Sekerka problem on the unit sphere from ? for a convergence experiment for our approximations as $\varepsilon\to 0$, see also ?.
@@
which means that the ODE system for $(\theta_{1},\theta_{2})$ from ? can be reduced to a scalar differential equation. The former is given by
@@
Combining (5.2.2) with (5.2) yields a scalar differential equation for $\theta_{1}$.
For the initial data we choose $R_{1}(0)=2R_{2}(0)=0.8$, so that $h_{1}(0)=-0.6$ and $h_{2}(0)=-\sqrt{0.84}$, respectively. In order to get the same time scale as in ?, we use $\alpha=\frac{\sqrt{2}}{3}$ and $\lambda=2$ for these experiments, and we visualize the phase field energies
@@
5.2.3 Spinodal decomposition on the unit sphere
In this section we consider $\mathcal{M}={\mathbb{S}}^{2}$ and let $\varepsilon=(16\pi)^{-1}$, as well as $N_{c}=1$, $N_{f}=128$, $\tau=10^{-6}$. We use different surface energy densities for evolutions that start from a random mixture with mean zero and values in $[-0.1,0.1]$. In this isotropic case, we observe the well known spinodal decomposition patterns, see Figure 6. When we use the anisotropy (5.1)(a), for which in the limit $\delta=0$ the normals $\nu=\pm e_{3}$ have the same energy density as the other six main facet normals of the crystalline Wulff shape, during the spinodal decomposition we can observe corners, see Figure 7. Note that the spatially homogeneous anisotropy leads to hexagonal symmetries only at the two poles, whereas the patterns near the equator resemble squares. The corresponding evolution for the anisotropy (5.1)(b) is shown in Figure 8. Here it can be seen that $\nu=\pm e_{3}$ is now the preferred normal direction, and so the interfaces very quickly align with it. As a final experiment, we also use the anisotropic density (5.1)(c), see Figure 9 for the numerical results. Now the opposite effect can be observed: since the normals $\nu=\pm e_{3}$ are relatively expensive, they are avoided by the developing interfaces. We also observe that out of the four spinodal decomposition experiments, only two settle on a global energy minimizer as described in Lemma 3.1. We conjecture that the final shape for the other two simulations is made up of “straight” segments that are aligned with the Wulff shape, in the sense that their normal vectors correspond to directions with minimal energy density. The anisotropic nature of the surface energy then seems to make it impossible to go from this local minimizer to one of the global ones.
@@ 5.2.4 Crystal growth on a sphere cap
=$\mathcal{M}=\{0\}\cup\left\{\begin{pmatrix}\tfrac{z_{1}}{|z|}\,\cos\theta(z)\\ \tfrac{z_{2}}{|z|}\,\cos\theta(z)\\ \sin\theta(z)\end{pmatrix}:z=\begin{pmatrix}z_{1}\\ z_{2}\\ 0\end{pmatrix}\in{\mathbb{R}}^{3},0<|z|\leq 1\right\},\ \theta(z)=(\tfrac{\pi}{2}+\tfrac{\pi}{18})(1-|z|)-\tfrac{\pi}{18}.$
For the physical parameters, similarly to ?, Fig. 9, we choose $\alpha=0.03$, $\rho=0.01$ and $w_{D}=-8$.
Some initial computations are shown in Figure 10, where for $\varepsilon=(32\pi)^{-1}$ we select the discretization parameters $N_{c}=16$, $N_{f}=128$ and $\tau=10^{-5}$. We start with an initial seed of radius $r_{0}=0.02$: either at the north pole, or on the equator on the $x$-axis, or on the equator on the $y$-axis. The evolutions in Figure 10 make it clear that the hexagonal aspect of the anisotropy (5.1)(a) only comes to the fore a the north pole. At the equator, on the other hand, we see interfaces with a fourfold symmetry, consistent with the Wulff shape displayed in Figure 3. We remark that these computations are for (4.3)(ii), compare with ?, (2.11)(ii). Repeating the same simulations for (4.3)(i) leads to the creation of boundary layers, which is why we prefer the choice (4.3)(ii) here and from now. We refer to ? for a more detailed discussion of this aspect.
Figure 10: ($\varepsilon=(32\pi)^{-1}$) Parameters as in ?, Fig. 9, but here $\varepsilon=(32\pi)^{-1}$ and $w_{D}=-8$. Starting seed on top (top), at the front (middle) and on the right (bottom). Displayed times are $t=0,0.05,0.1$ (top) and $t=0,0.01,0.02$ (middle and bottom).
Having established that for the homogeneous hexagonal anisotropies in this section it is beneficial for sixfold structures to start with a seed at the north pole, we present a few more computations with a seed of radius $r_{0}=0.02$ at the north pole. Then for different physical parameters, we obtain slightly different evolutions. But note that what they all have in common is that the crystal will initially grow sixth arms, but these will then often display non-hexagonal sidearms and structures. The simulations shown in Figure 11 are for $w_{D}=-12$, with either $\rho=\alpha=0.01$, $\rho=10\alpha=0.01$ or $\rho=\alpha=0.001$, respectively. Here we let $\varepsilon=(32\pi)^{-1}$ and $N_{c}=32$, $N_{f}=256$ and $\tau=10^{-5}$.
@@ 5.3 Consistent 2d anisotropies on the unit sphere
In this subsection we present some numerical evidence in support of the construction of consistent anisotropies along a surface as proposed in §3.2. We first repeat the simulations in Figure 10, but now for the anisotropy (3.2) with (3.7), for $L=3$, the usual 2d hexagonal anisotropy from, e.g., ????, i.e.,
@@
When we repeat the simulations in Figure 12 for the new anisotropy these mushy regions disappear, see Figure 16. Observe that compared to the dendritic growth in the former case, the new simulations show the natural six-fold growth also way beyond the equator. In fact, the interface growths in Figure 17 look very similar to the 2d simulations in ???.

### v2

@@ 4 Finite element approximation
Let $\mathcal{M}^{h}$ be a polyhedral hypersurface approximating $\mathcal{M}$, and let $\{{\cal T}^{h}\}_{h>0}$ be a family of open triangles with $\mathcal{M}^{h}=\cup_{\sigma\in{\cal T}^{h}}\overline{\sigma}$. We refer to [16] for more details on polyhedral approximations of surfaces and finite element spaces on polyhedral surfaces. Associated with ${\cal T}^{h}$ is the finite element space
@@ 4.1 The obstacle potential
It is possible to generalize the considered model in the following ways. All of the results presented in this paper remain valid for these generalizations, on using the techniques developed by the authors in [12].
@@
    can be considered, so that (3.4) collapses to the case $r=1$. In the case $r>1$, the term $B(\cdot,\nabla_{\!s}\Phi^{n-1})$ in, e.g., (4.2b) needs to be changed to $B_{r}(\cdot,\nabla_{\!s}\Phi^{n-1},\nabla_{\!s}\Phi^{n})$, where the matrices $B_{r}$ are defined analogously to the flat and spatially homogeneous case in [12].
@@
Proof. The existence and uniqueness results follow analogously to Lemma 3.1 in [12], and so we omit their proof here.
@@ 4.2 Smooth potentials
and for simplicity we restrict our attention to that case. Details on how to deal with more general potentials can be found in [12].
@@
Proof. The existence and uniqueness results follow analogously to Theorem 3.8 in [12]. The proof of the stability bound (4.2) is similar to the proof of Theorem 4.2, making use of the convex/concave splitting $\Psi^{\prime}=\phi=\phi^{+}+\phi^{-}$.
@@ 5 Numerical results
We implemented the scheme (4.2) with the help of the finite element toolbox ALBERTA, see [55]. To increase computational efficiency, we employ adaptive meshes, which have a finer mesh size $h_{f}\approx\frac{1}{N_{f}}$ within the diffuse interfacial regions and a coarser mesh size $h_{c}\approx\frac{1}{N_{c}}$ away from them, with $N_{f},N_{c}\in{\mathbb{N}}$, see [18, 17] for a more detailed description in the planar case. For the solution of the nonlinear systems of equations we employ the iterative solvers discussed in [12].
@@ 5.1 Spatially inhomogeneous anisotropies in 2d
Then, on $\mathcal{M}=(-\frac{3}{2},\frac{3}{2})\times(-\frac{1}{2},\frac{1}{2})\times\{0\}$ we use the spatially weighted isotropic surface energy
@@
Figure 2: ($\varepsilon=(16\pi)^{-1}$) Solutions for the Cahn–Hilliard equation in the case of a spatially inhomogeneous energy in the plane. Snapshots of the evolutions at times $t=0,0.1,0.2,0.3,0.4,0.5,0.6,0.7,2$.
@@ 5.2 Spatially homogeneous anisotropies in 3d
where $R_{1}(\theta)=\left(\!\!\!\scriptsize\begin{array}[]{rrr}\cos\theta&\sin\theta&0\\ -\sin\theta&\cos\theta&0\\ 0&0&1\end{array}\!\!\right)$, $R_{2}(\theta)=\left(\!\!\!\scriptsize\begin{array}[]{rrr}\cos\theta&0&\sin\theta\\ 0&1&0\\ -\sin\theta&0&\cos\theta\end{array}\!\!\right)$ and $l_{\delta}({p})=\left[\delta^{2}\,|{p}|^{2}+p_{1}^{2}\,(1-\delta^{2})\right]^{\frac{1}{2}}$. The Wulff shapes for (5.1) are shown in Figure 3. We recall from [36] that for a given anisotropy $\gamma_{0}$ the boundary of its Wulff shape is the solution of the isoperimetric problem for $\mathcal{E}_{0}(\Gamma)=\int_{\Gamma}\gamma_{0}(\nu)\;{\rm d}{\mathcal{H}}^{2}$. Hence they are often used to visualize the properties of $\gamma_{0}$.
@@
We note that for $\delta=0$, the densities (5.1) represent crystalline anisotropies, whose Wulff shapes are given by hexagonal prisms, cf. [9, Fig. 3]. The coefficients in (5.1)(a) are chosen such that, in the limit $\delta=0$, the eight admissible normal directions all have the same energy, while in (5.1)(b) and (5.1)(c) the two normals ±$\pm e_{3}$ have a smaller/larger energy than the other six directions.
We use the anisotropy (5.1)(b) with $\delta=10^{-2}$ to numerically investigate the statement from Lemma 3.1. To this end, we choose as initial interface a perturbed great circle on the unit sphere, that lies in a hyperplane that makes a certain fixed angle $\theta$ with the $x-z$-plane. In particular, in Figure 4 we show simulations for $\theta=10^{\circ},30^{\circ},60^{\circ}$. Here we let $\varepsilon=(16\pi)^{-1}$, $N_{c}=1$, $N_{f}=128$ and $\tau=10^{-4}$. We note that each simulation settles on a great circle on the sphere. Here we note that as $\nu_{\min}=e_{3}$, the equator has the lowest energy of all the great circles, and so if the initial curve on $\mathcal{M}$ is sufficiently inclined, that global minimizer is indeed reached by the evolution. For only small deviations from a north-south great circle, the evolutions settle on that local minimizer instead. We remark that the discrete steady state solutions exhibit final discrete energies $\mathcal{F}^{h}(W^{n},\Phi^{n})$ of 12.7, 6.7 and 6.7, respectively. This once again confirms that the equator has the lowest energy.
Figure 4: ($\varepsilon=(16\pi)^{-1}$) Anisotropic Cahn–Hilliard equation on a sphere. Snapshots of the evolutions at times $t=0,1,5$ (from left to right). The initial data approximates a great circle within a hyperplane that is tilted by $10^{\circ},30^{\circ},60^{\circ}$ (from top to bottom).
@@ 5.2.2 Convergence experiment on the unit sphere
We use the rotationally symmetric solution for the Mullins–Sekerka problem on the unit sphere from [52] for a convergence experiment for our approximations as $\varepsilon\to 0$, see also [7].
@@
which means that the ODE system for $(\theta_{1},\theta_{2})$ from [52] can be reduced to a scalar differential equation. The former is given by
@@
Combining (5.3) with (5.2) yields a scalar differential equation for $\theta_{1}$.
For the initial data we choose $R_{1}(0)=2R_{2}(0)=0.8$, so that $h_{1}(0)=-0.6$ and $h_{2}(0)=-\sqrt{0.84}$, respectively. In order to get the same time scale as in [52], we use $\alpha=\frac{\sqrt{2}}{3}$ and $\lambda=2$ for these experiments, and we visualize the phase field energies
@@
5.2.3 Spinodal decomposition on closed surfaces
In this section we consider $\varepsilon=(16\pi)^{-1}$, as well as $N_{c}=1$, $N_{f}=128$, $\tau=10^{-6}$. We use different surface energy densities for evolutions that start from a random mixture with mean zero and values in $[-0.1,0.1]$. At first we choose the unit sphere $\mathcal{M}={\mathbb{S}}^{2}$. In this isotropic case, we observe the well known spinodal decomposition patterns, see Figure 6. When we use the anisotropy (5.1)(a), for which in the limit $\delta=0$ the normals $\nu=\pm e_{3}$ have the same energy density as the other six main facet normals of the crystalline Wulff shape, during the spinodal decomposition we can observe corners, see Figure 7. Note that the spatially homogeneous anisotropy leads to hexagonal symmetries only at the two poles, whereas the patterns near the equator resemble squares. The corresponding evolution for the anisotropy (5.1)(b) is shown in Figure 8. Here it can be seen that $\nu=\pm e_{3}$ is now the preferred normal direction, and so the interfaces very quickly align with it. As a final experiment, we also use the anisotropic density (5.1)(c), see Figure 9 for the numerical results. Now the opposite effect can be observed: since the normals $\nu=\pm e_{3}$ are relatively expensive, they are avoided by the developing interfaces. We also observe that out of the four spinodal decomposition experiments, only two settle on a global energy minimizer as described in Lemma 3.1. We conjecture that the final shape for the other two simulations is made up of “straight” segments that are aligned with the Wulff shape, in the sense that their normal vectors correspond to directions with minimal energy density. The anisotropic nature of the surface energy then seems to make it impossible to go from this local minimizer to one of the global ones.
@@ 5.2.3 Spinodal decomposition on closed surfaces
We end this subsection with a simulation on a surface that is the boundary of a nonconvex domain. In particular, we base $\mathcal{M}^{h}$ on a rescaled version of the biconcave disk obtained in [13, Fig. 8]. The total dimensions of $\mathcal{M}^{h}$ are about $0.28\times 1\times 1$. On this surface we repeat the simulation from Figure 7. The results of this new numerical experiment are shown in Figure 10, where we once again observe the squared patterns due to the alignment of the surface with respect to the Wulff shape of the anisotropy (5.1)(a).
Figure 10: ($\varepsilon=(16\pi)^{-1}$) Spinodal decomposition on a biconcave disk, for (5.1)(a). Snapshots of the evolutions at times $t=10^{-4},2\times 10^{-4},5\times 10^{-4},10^{-3},5\times 10^{-3},0.01$.
@@ 5.2.4 Crystal growth on a sphere cap
=$\mathcal{M}=\left\{\begin{pmatrix}0\\ 0\\ 1\end{pmatrix}\right\}\cup\left\{\begin{pmatrix}\tfrac{z_{1}}{|z|}\,\cos\theta(z)\\ \tfrac{z_{2}}{|z|}\,\cos\theta(z)\\ \sin\theta(z)\end{pmatrix}:z=\begin{pmatrix}z_{1}\\ z_{2}\\ 0\end{pmatrix}\in{\mathbb{R}}^{3},0<|z|\leq 1\right\},$
$\text{where}\ \theta(z)=(\tfrac{\pi}{2}+\tfrac{\pi}{18})(1-|z|)-\tfrac{\pi}{18}.$
For the physical parameters, similarly to [12, Fig. 9], we choose $\alpha=0.03$, $\rho=0.01$ and $w_{D}=-8$.
Some first computations are shown in Figure 11, where for $\varepsilon=(32\pi)^{-1}$ we select the discretization parameters $N_{c}=16$, $N_{f}=128$ and $\tau=10^{-5}$. We start with an initial seed of radius $r_{0}=0.02$: either at the north pole, or on the equator on the $x$-axis, or on the equator on the $y$-axis. The evolutions in Figure 11 make it clear that the hexagonal aspect of the anisotropy (5.1)(a) only comes to the fore at the north pole. At the equator, on the other hand, we see interfaces with a four-fold symmetry, consistent with the Wulff shape displayed in Figure 3. We remark that these computations are for (4.3)(ii), compare with [12, (2.11)(ii)]. Repeating the same simulations for (4.3)(i) leads to the creation of boundary layers, which is why we prefer the choice (4.3)(ii) here and from now. We refer to [12] for a more detailed discussion of this aspect.
Figure 11: ($\varepsilon=(32\pi)^{-1}$) Parameters as in [12, Fig. 9], but here $\varepsilon=(32\pi)^{-1}$ and $w_{D}=-8$. Starting seed on top (top), at the front (middle) and on the right (bottom). Displayed times are $t=0,0.05,0.1$ (top) and $t=0,0.01,0.02$ (middle and bottom).
Having established that for the homogeneous hexagonal anisotropies in this section it is beneficial for six-fold structures to start with a seed at the north pole, we present a few more computations with a seed of radius $r_{0}=0.02$ at the north pole. Then for different physical parameters, we obtain slightly different evolutions. But note that what they all have in common is that the crystal will initially grow sixth arms, but these will then often display non-hexagonal sidearms and structures. The simulations shown in Figure 12 are for $w_{D}=-12$, with either $\rho=\alpha=0.01$, $\rho=10\alpha=0.01$ or $\rho=\alpha=0.001$, respectively. Here we let $\varepsilon=(32\pi)^{-1}$ and $N_{c}=32$, $N_{f}=256$ and $\tau=10^{-5}$.
@@
We stress once more that if the seed would not be placed at the north pole, or if the ambient anisotropy would be rotated so that the tangent space at the north pole is not aligned with the six-fold symmetry, then the resulting growth would not show the desired six-fold symmetric patterns.
@@ 5.3 Consistent 2d anisotropies on the unit sphere
In this subsection we present some numerical evidence in support of the construction of consistent anisotropies along a surface as proposed in §3.2. We first repeat the simulations in Figure 11, but now for the anisotropy (3.2) with (3.7), for $L=3$, the usual 2d hexagonal anisotropy from, e.g., [4, 7, 10, 12], i.e.,
@@
When we repeat the simulations in Figure 13 for the new anisotropy these mushy regions disappear, see Figure 17. Observe that compared to the dendritic growth in the former case, the new simulations show the natural six-fold growth also way beyond the equator. In fact, the interface growths in Figure 18 look very similar to the 2d simulations in [7, 8, 9].